# Xinyi Bai2, Weijie J. Su1, Camillo J. Taylor1, Tanwi Mallick3

### v4

@@ Abstract
Rationality is the quality of being guided by reason, characterized by decision-making that aligns with evidence and logical principles. It plays a crucial role in reliable problem-solving by ensuring well-grounded and consistent solutions. While large language models (LLMs) have made significant progress in generating human-like text, they still exhibit limitations such as bounded knowledge space and inconsistent outputs. In response, recent efforts have shifted toward developing multimodal and multi-agent systems, as well as integrating modules like external tools, programming codes, symbolic reasoners, utility function, and conformal risk controls rather than relying solely on a single LLM for decision-making. This paper surveys the state-of-the-art advancements in language and multimodal agents, evaluates how they contribute to make intelligent agents more rational, and identifies open challenges and future research directions. We maintain an open repository at https://github.com/bowen-upenn/Agent_Rationality.
@@ 1 Introduction
To address these challenges, research is shifting toward multimodal agents and multi-agent systems, as complex problems in real life often require collaborations of experts across fields Eisenführ et al. (2010) and data from diverse sources. Formally, we refer "agent" Bommasani et al. (2021); Park et al. (2023) as an artificial intelligent entity that perceives and understands its environment through various inputs — either natural language or multimodal information like vision, audio, and programming codes — and acts to achieve specific goals or tasks within natural language domains. This can encompass a range of systems, from a single LLM, a multimodal foundation model with instruction following capabilities Liu et al. (2024b, 2023b), to a multi-agent system that integrates multiple AI agents, traditional machine learning or symbolic reasoning modules, external knowledge bases, and tools working together towards a collective goal within the same environment.
@@ 2 Scope
Existing surveys in the field of agent systems, such as those on multimodal Xie et al. (2024a); Durante et al. (2024); Cui et al. (2024); Xu et al. (2024d); Li et al. (2024a) and multi-agent systems Han et al. (2024); Guo et al. (2024); Zhang et al. (2024b); Cheng et al. (2024), primarily focus on their components, architectures, profiling, planning, communication strategies, memory mechanisms, and applications. Additionally, many works Jiang et al. (2024a); Wei et al. (2022); Yao et al. (2024); Valmeekam et al. (2023); Xu et al. (2024a); Prasad et al. (2023); Khardon and Roth (1997); Huang and Chang (2022); Zhang et al. (2024b); Qiao et al. (2022) and surveys Qiao et al. (2022); Huang and Chang (2022); Ahn et al. (2024); Giadikiaroglou et al. (2024); Liang et al. (2024); Zheng et al. (2024c); Zhang et al. (2024b) explore the reasoning capabilities of LLMs. Although reasoning plays an important role in ensuring rationality, especially in complex scenarios, it remains parallel to our focus, as mentioned in Section 1. Furthermore, some works touch the aspect of rationality in LLMs Kassner et al. (2023); Raman et al. (2024); Macmillan-Scott and Musolesi (2024), but they focus on one specific algorithm or application domain. To the best of our knowledge, this survey is the first to comprehensively explore the notion of rationality in language and multimodal agents. We aim to bridge the gap between rationality and agent system, analyzing how designs in these agents and agent systems contribute to advancing certain key axioms of rationality.
@@ 3 Defining Rationality in Agents
A rational agent’s decision-making should be grounded in physical and factual reality, incorporating information it perceives from multimodal formats and sources.
@@ 4.1.1 Grounding on multimodal information
Grounding an agent solely based on language can be challenging. As a picture is worth a thousand words, recent advances in large multimodal models Li et al. (2024a) integrate language, vision, and other sensory modalities to offer a more comprehensive grounding of information, thereby enhancing the understanding of decision-making contexts. Multimodal foundation models, including but not limited to CLIP Radford et al. (2021), VLBERT and ViLBERT Su et al. (2019); Lu et al. (2019), BLIP-2 Li et al. (2023d), UniAudio Yang et al. (2023a), AudioLM Borsos et al. (2023), TANGO Ghosal et al. (2023), SpeechGPT Zhang et al. (2023), (Open) Flamingo Alayrac et al. (2022); Awadalla et al. (2023), LLaVA Liu et al. (2024b, 2023b), CogVLM Wang et al. (2023d), MiniGPT-4 Zhu et al. (2023), Whisper Radford et al. (2023), GPT-4 Vision OpenAI (2023) and GPT-4o OpenAI (2024), LLaMA 3.2 Meta (2024), and Gemini 1.5 Pro Reid et al. (2024) serve as the cornerstones for downstream tasks in multimodal agent systems.
The adaptation of Reinforcement Learning from Human Feedback (RLHF) OpenAI (2024); Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022); Zhang et al. (2024a), a technique popularized in language-only models, also demonstrates promising advancements in reducing hallucination from cross-modal misalignment Sun et al. (2023). Visual instruction-tuning Liu et al. (2024b); Dai et al. (2024); Bai et al. (2023); Wang et al. (2023b) also enables foundation models to engage in more detailed multi-round, context-aware human-agent interactions and collaborations with other agents. This opens the possibility of the System 2 process Kahneman (2011) in multimodal models.
Multi-modalities help expand the functionality of agents by allowing them to access more comprehensive and diverse data. For example, Chain-of-Action Pan et al. (2024) advances the single-modal Search-in-the-Chain Xu et al. (2023) by supporting multimodal data retrieval for faithful question answering. multimodal understanding in DoraemonGPT Yang et al. (2024) is necessary for spatial-temporal videos analysis. LogicVista  expands logical reasoning capabilities to visual contexts. The multimodal capabilities also allow HuggingGPT Shen et al. (2024d), Agent LUMOS Yin et al. (2023), ToolAlpaca Tang et al. (2023), and AssistGPT Gao et al. (2023a) to expand the scope of tasks they can address, including cooperation among specialized agents or tools capable of handling different modalities. Web agents like Zheng et al. (2024a); Shen et al. (2024c); Deng et al. (2024); Gur et al. (2023) grounded on the graphical user interface (GUI) offers higher information density compared to solely HTML codes in textual formats Yao et al. (2022); Nakano et al. (2021).
@@ 4.2.1 Consensus from reflection and multi-agent collaboration
Multi-round self-reflection that encourages a single LLM to critically evaluate its previous responses has demonstrated the effectiveness of deliberation in improving logical coherence and qualities OpenAI (2024); Shinn et al. (2024); Madaan et al. (2024); Wang et al. (2022b); Zhong et al. (2024); Lu et al. (2023); Xu et al. (2024c, a); Xiong et al. (2024b). Expanding on this, multi-agent systems introduce collaboration among multiple agents Zhang et al. (2024c), enabling collective deliberation through cross-examination and debate. For instance, LM vs LM Cohen et al. (2023) introduces a cross-examination between two agents to detect errors and make factuality decisions. FORD Xiong et al. (2023a) reports an accuracy increase up to $4.9\%$ compared to a single LLM. AgentReview Jin et al. (2024) presents how discussions cause distribution shifts in final reviews compared to initial reviews. Liang et al. (2023) demonstrates the superiority of multi-agent debate over self-reflection, with final consensus achieving a $16.0\%$ improvement in reasoning tasks. Similarly, Du et al. (2023) finds that LLMs can converge on a single shared answer after multiple rounds of debate, resulting in a factual accuracy increase of $7.2$-$15.9\%$ across tasks. All these approaches enhance the system’s capability to capture initial errors, improve factuality in reasoning, and achieve final consensus with fewer inconsistencies.
@@ 4.3.1 Representation invariance across modalities
The concept of invariance is the cornerstone of Visual Question Answering (VQA) agents Chen et al. (2022); Jiang et al. (2024b); Wang et al. (2023e); Yi et al. (2018); Wang et al. (2022a); Bao et al. (2022); Li et al. (2024c); Zhao and Xu (2023). On one hand, these agents must grasp the invariant semantics of any open-ended questions posed about images, maintaining consistency despite variations in wording, syntax, or language. On the other hand, within a multi-agent VQA system, visual agents can provide crucial verification and support for language-based reasoning Wang et al. (2023e); Jiang et al. (2024b); Zhao and Xu (2023), while language queries can direct the attention of visual agents, based on a shared and invariant underlying knowledge across vision and language domains.
@@ 5.3 Evaluating Invariance from Irrelevant Information
Studies such as Shi et al. (2023), Wu et al. (2024), Liu et al. (2024c), and Yoran et al. (2023) investigate the phenomenon of “lost-in-context" by introducing random or misleading sentences into original problem statements. Early benchmarks such as those by Weston et al. (2015), Sinha et al. (2019), Clark et al. (2020), and Webson and Pavlick (2021) also incorporate irrelevant content. More recent benchmarks like MileBench Song et al. (2024), Mementos Wang et al. (2024c), Seed-bench-2 Li et al. (2023b), and DEMON Li et al. (2023c) extend these evaluations to multimodal agents acting in long context with image sequences. In these scenarios, the agents must accurately isolate relevant information from large context windows.
@@ 5.4 Evaluating Orderability of Preference
Having an orderability of preference is essential when leveraging LLMs as evaluators to ensure reliable assessments. Luo et al. (2023); Shen et al. (2023); Gao et al. (2023b); Wang et al. (2023a); Chen et al. (2023); Chiang and Lee (2023b); Zheng et al. (2024b); Fu et al. (2023); Liu et al. (2023d) highlight challenges with LLMs in this role, reporting inconsistent ratings and difficulties in establishing reliable comparisons. This inconsistencies raise concerns about the ability of LLMs to accurately rank and evaluate different options or responses. ChatEval Chan et al. (2023) and Bai et al. (2024) suggest improved preference aligned with humans through multi-agent collaborations. The multiple choice problems PaperswithcodeMCQA serves as another common testing ground. (Zong et al., 2023; Zheng et al., 2023a) show that LLMs are susceptible to the rearranging of options, often failing to maintain a coherent order of preference.
@@ 6 Open Problems and Future Directions
Current methods are neither sufficient nor necessary to achieve human-level rationality, but they serve as instrumental tools that bridge the gap between an LLM’s response and rationality. These approaches enable agent systems to more closely mimic rational thinking in their output responses from the end-user’s perspective. However, how to effectively close the loop and bake these more rational outputs back into foundation models themselves Zhao et al. (2024) beyond mere fine-tuning remains an open question. In other words, can we leverage these more rational outputs to inherently enhance a single foundation model’s rationality in its initial responses?
@@ 7 Discussion
Rationality adds another crucial dimension to the performance of intelligent agents besides reasoning capabilities. It becomes increasingly important for human users applying these agents in critical sectors like health care and finance that expect consistent and reliable decision-making. Meanwhile, more rational agents - capable of providing coherent orderability of preferences and grounded in knowledge beyond their bounded parametric memory - could enhance their roles in automatic evaluation and preference alignment, where human involvements are expensive yet unstable.

### v6

@@ Abstract
This work discusses how to build more rational language and multimodal agents and what criteria define rationality in intelligent systems. Rationality is the quality of being guided by reason, characterized by decision-making that aligns with evidence and logical principles. It plays a crucial role in reliable problem-solving by ensuring well-grounded and consistent solutions. Despite their progress, large language models (LLMs) often fall short of rationality due to their bounded knowledge space and inconsistent outputs. In response, recent efforts have shifted toward developing multimodal and multi-agent systems, as well as integrating modules like external tools, programming codes, symbolic reasoners, utility function, and conformal risk controls rather than relying solely on a single LLM for decision-making. This paper surveys state-of-the-art advancements in language and multimodal agents, assesses their role in enhancing rationality, and outlines open challenges and future research directions. We maintain an open repository at https://github.com/bowen-upenn/Agent_Rationality.
@@ 1 Introduction
To address these challenges, research is shifting toward multimodal agents and multi-agent systems, as complex problems in real life often require collaborations of experts across fields Eisenführ et al. (2010) and data from diverse sources. Formally, we refer "agent" Bommasani et al. (2021) as an artificial intelligent entity that perceives and understands its environment through various inputs — either natural language or multimodal information like vision, audio, and codes — and acts to achieve specific goals or tasks within natural language domains, while the term "agentic" to describe such behaviors Kapoor et al. (2024); Ng (2024). This can encompass a range of systems, from a single LLM, a multimodal foundation model with instruction following capabilities Liu et al. (2024b, 2023b), to a multi-agent system that integrates multiple AI agents, traditional machine learning or symbolic reasoning modules, external knowledge bases, and tools working together towards a collective goal within the same environment.
@@ 2 Scope
Existing surveys in the field of agent systems, such as those on multimodal Xie et al. (2024a); Durante et al. (2024); Cui et al. (2024); Xu et al. (2024d); Li et al. (2024a) and multi-agent systems Han et al. (2024); Guo et al. (2024); Zhang et al. (2024c); Cheng et al. (2024), primarily focus on their components, architectures, profiling, planning, communication strategies, memory mechanisms, and applications. Additionally, many works Jiang et al. (2024a); Wei et al. (2022); Yao et al. (2024); Cai et al. (2024); Valmeekam et al. (2023); Xu et al. (2024a); Prasad et al. (2023); Khardon and Roth (1997); Huang and Chang (2022); Zhang et al. (2024c); Qiao et al. (2022) and surveys Qiao et al. (2022); Huang and Chang (2022); Ahn et al. (2024); Giadikiaroglou et al. (2024); Liang et al. (2024); Zheng et al. (2024c); Zhang et al. (2024c); Xiong et al. (2024b) explore the reasoning capabilities of LLMs. Although reasoning plays an important role in ensuring rationality, especially in complex scenarios, it remains parallel to our focus, as mentioned in Section 1. Furthermore, some works touch the aspect of rationality in LLMs Kassner et al. (2023); Raman et al. (2024); Macmillan-Scott and Musolesi (2024), but they focus on one specific algorithm or application domain. To the best of our knowledge, this survey is the first to comprehensively explore the notion of rationality in language and multimodal agents. We aim to bridge the gap between rationality and agent system, analyzing how designs in these agents and agent systems contribute to advancing certain key axioms of rationality.
@@ 3 Defining Rationality in Agents
A rational agent’s decision-making should be grounded in physical and factual reality, incorporating information it perceives from multimodal formats and sources. In contrast, an irrational agent generates hallucinations Huang et al. (2023), producing false or misleading information that is not grounded in facts.
@@ 4.1.1 Grounding on multimodal information
Grounding an agent solely based on language can be challenging. As a picture is worth a thousand words, recent advances in large multimodal models Li et al. (2024a) integrate language, vision, and other sensory modalities to offer a more comprehensive grounding of information, thereby enhancing the understanding of decision-making contexts. Multimodal foundation models, including but not limited to CLIP Radford et al. (2021), VLBERT and ViLBERT Su et al. (2019); Lu et al. (2019), BLIP-2 Li et al. (2023d), UniAudio Yang et al. (2023a), AudioLM Borsos et al. (2023), TANGO Ghosal et al. (2023), SpeechGPT Zhang et al. (2023), (Open) Flamingo Alayrac et al. (2022); Awadalla et al. (2023), LLaVA Liu et al. (2024b, 2023b), CogVLM Wang et al. (2023d), MiniGPT-4 Zhu et al. (2023), Whisper Radford et al. (2023), GPT-4 Vision OpenAI (2023) and GPT-4o OpenAI (2024), LLaMA 3.2 Meta (2024), and Gemini 1.5 Pro Reid et al. (2024) serve as the cornerstones for downstream tasks in multimodal agent systems. More agentic systems increasingly depend on multimodal information to enhance multimodal reasoning. Zhang et al. (2024a); Brienza et al. (2024); Elhenawy et al. (2024); Chen et al. (2023a); Dong et al. (2024); Wu et al. (2024b).
The adaptation of Reinforcement Learning from Human Feedback (RLHF) OpenAI (2024); Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022); Zhang et al. (2024b), a technique popularized in language-only models, also demonstrates promising advancements in reducing hallucination from cross-modal misalignment Sun et al. (2023). Visual instruction-tuning Liu et al. (2024b); Dai et al. (2024); Bai et al. (2023); Wang et al. (2023b) also enables foundation models to engage in more detailed multi-round, context-aware human-agent interactions and collaborations with other agents. This opens the possibility of the System 2 process Kahneman (2011) in multimodal models.
Multi-modalities help expand the functionality of agents by allowing them to access more comprehensive and diverse data. For example, Chain-of-Action Pan et al. (2024) advances the single-modal Search-in-the-Chain Xu et al. (2023) by supporting multimodal data retrieval for faithful question answering. multimodal understanding in DoraemonGPT Yang et al. (2024) is necessary for spatial-temporal videos analysis. LogicVista Xiao et al. (2024) expands logical reasoning capabilities to visual contexts. The multimodal capabilities also allow HuggingGPT Shen et al. (2024d), Agent LUMOS Yin et al. (2023), ToolAlpaca Tang et al. (2023), and AssistGPT Gao et al. (2023a) to expand the scope of tasks they can address, including cooperation among specialized agents or tools capable of handling different modalities. Web agents like Zheng et al. (2024a); Shen et al. (2024c); Deng et al. (2024); Gur et al. (2023); Zhou et al. (2023); Koh et al. (2024) grounded on the graphical user interface (GUI) offers higher information density compared to solely HTML codes in textual formats Yao et al. (2022); Nakano et al. (2021).
@@ 4.2.1 Consensus from reflection and multi-agent collaboration
Multi-round self-reflection that encourages a single LLM to critically evaluate its previous responses has demonstrated the effectiveness of deliberation in improving logical coherence and qualities OpenAI (2024); Shinn et al. (2024); Madaan et al. (2024); Wang et al. (2022b); Zhong et al. (2024); Lu et al. (2023); Xu et al. (2024c, a); Xiong et al. (2024c). Expanding on this, multi-agent systems introduce collaboration among multiple agents Zhang et al. (2024d), enabling collective deliberation through cross-examination and debate. For instance, LM vs LM Cohen et al. (2023) introduces a cross-examination between two agents to detect errors and make factuality decisions. FORD Xiong et al. (2023a) reports an accuracy increase up to $4.9\%$ compared to a single LLM. AgentReview Jin et al. (2024) presents how discussions cause distribution shifts in final reviews compared to initial reviews. Liang et al. (2023) demonstrates the superiority of multi-agent debate over self-reflection, with final consensus achieving a $16.0\%$ improvement in reasoning tasks. Zhao et al. (2023) investigates the multi-agent competition behaviors. Similarly, Du et al. (2023) finds that LLMs can converge on a single shared answer after multiple rounds of debate, resulting in a factual accuracy increase of $7.2$-$15.9\%$ across tasks. All these approaches enhance the system’s capability to capture initial errors, improve factuality in reasoning, and achieve final consensus with fewer inconsistencies.
@@ 4.3.1 Representation invariance across modalities
The concept of invariance is the cornerstone of Visual Question Answering (VQA) agents Chen et al. (2022); Jiang et al. (2024b); Wang et al. (2023e); Yi et al. (2018); Wang et al. (2022a); Bao et al. (2022); Li et al. (2024c); Zhao and Xu (2023); Fan et al. (2024). On one hand, these agents must grasp the invariant semantics of any open-ended questions posed about images, maintaining consistency despite variations in wording, syntax, or language. On the other hand, within a multi-agent VQA system, visual agents can provide crucial verification and support for language-based reasoning Wang et al. (2023e); Jiang et al. (2024b); Zhao and Xu (2023), while language queries can direct the attention of visual agents, based on a shared and invariant underlying knowledge across vision and language domains.
@@ 5.3 Evaluating Invariance from Irrelevant Information
Studies such as Shi et al. (2023), Wu et al. (2024a), Liu et al. (2024c), and Yoran et al. (2023) investigate the phenomenon of “lost-in-context" by introducing random or misleading sentences into original problem statements. Early benchmarks such as those by Weston et al. (2015), Sinha et al. (2019), Clark et al. (2020), and Webson and Pavlick (2021) also incorporate irrelevant content. More recent benchmarks like MileBench Song et al. (2024), Mementos Wang et al. (2024c), Seed-bench-2 Li et al. (2023b), and DEMON Li et al. (2023c) extend these evaluations to multimodal agents acting in long context with image sequences. In these scenarios, the agents must accurately isolate relevant information from large context windows.
@@ 5.4 Evaluating Orderability of Preference
Having an orderability of preference is essential when leveraging LLMs as evaluators to ensure reliable assessments. Luo et al. (2023); Shen et al. (2023); Gao et al. (2023b); Wang et al. (2023a); Chen et al. (2023b); Chiang and Lee (2023b); Zheng et al. (2024b); Fu et al. (2023); Liu et al. (2023d) highlight challenges with LLMs in this role, reporting inconsistent ratings and difficulties in establishing reliable comparisons. This inconsistencies raise concerns about the ability of LLMs to accurately rank and evaluate different options or responses. ChatEval Chan et al. (2023) and Bai et al. (2024) suggest improved preference aligned with humans through multi-agent collaborations. The multiple choice problems PaperswithcodeMCQA serves as another common testing ground. (Zong et al., 2023; Zheng et al., 2023a) show that LLMs are susceptible to the rearranging of options, often failing to maintain a coherent order of preference.
@@ 6 Open Problems and Future Directions
Despite significant research efforts in this survey aimed at achieving one or more axioms of rationality, most existing algorithms rely on external tools, thereby not inherently enhancing the rationality of artificial intelligence. In other words, current methods are neither sufficient nor necessary to achieve human-level rationality, but they serve as instrumental tools that bridge the gap between an LLM’s response and rationality. These approaches enable agent systems to more closely mimic rational thinking in their output responses from the end-user’s perspective. However, how to effectively close the loop and bake these more rational outputs back into foundation models themselves Zhao et al. (2024) beyond mere fine-tuning remains an open question. It remains a question if we can leverage these more rational outputs, or training the model to verify against rational axioms for reward scores, to inherently enhance a single foundation model’s rationality in its initial responses without external assistance.
@@ 7 Discussion
Rationality adds another crucial dimension to the performance of intelligent agents besides reasoning capabilities. More rational agents - capable of providing coherent orderability of preferences and grounded in knowledge beyond their bounded parametric memory - could enhance their roles in automatic evaluation and preference alignment, where human involvements are expensive yet unstable.
It becomes increasingly important for human users applying these agents in critical sectors like health care and finance that expect consistent and reliable decision-making. For example, in financial domains, decision-making by LLMs must operate within an acceptable risk threshold. By adapting conformal risks, we can enhance the orderability of preference and create a theoretical framework for parameterizing risk preferences into utility models. This approach enables rational financial agents to optimize decisions to max utilities while controlling risks through loss functions that bound expected risks.
Medical domains also present unique multi-modal challenges, particularly in medical imaging interpretation. For instance, diagnostic methods must withstand irrelevant information like background noise in MRI scans. An VLM could provide diagnostic explanations by looking at these images, and an LLM could collaborate with it as a multi-agent system to verify against pre-defined illness criteria, incorporating external knowledge retrieval to improve information grounding. Besides, a multi-agent collaboration could involve diverse medical-domain personas, such as surgeons, nurses, physicians, radiologists, pharmacists, offering diagnostic recommendations from different perspectives, potentially reducing hallucinations through collaborations. As a result, we hope our survey could serve as a tool-box for agent developers who want to build more rational agents in diverse domains.
@@ 8 Limitations
9 Acknowledgement
This work was funded by the Laboratory Directed Research and Development (LDRD) program at Argonne National Laboratory, with support from the Office of Science, U.S. Department of Energy, under Contract No. DEAC02-06CH11357. This work was also supported by NSF grant CCF-2112665 (TILOS), which provides funding for B. Jiang. and C. J. Taylor. Y. Xie and W. J. Su also acknowledge support from the NSF HDR TRIPODS award (CCF-1934876).